# NORDITA 2025-061

### v1

@@ I Introduction
As mentioned above, one well-known decay mechanism of local cosmic strings involves the nucleation of monopole–antimonopole pairs along the string core [6], a process whose GW signatures were discussed for example in [49, 50, 51, 52, 53, 54] (see [43, 55, 56] for an analysis in the context of the latest NANOGrav data). In contrast, our scenario introduces a qualitatively new decay channel: instead of fragmenting through monopole pair production, the string network decays via $O(2)\times O(2)$ bubble nucleation. This process leads to unique features in the GW spectrum, making it potentially distinguishable from the monopole-induced scenario.
While a complete study of these signatures is left for future work, we mention the possibility that such a combined signal could be tested against the GW background observed recently by PTAs. To make this more concrete, we introduce a simple two-field model that generates a cosmic-string network with sizeable tension and subsequently triggers its delayed decay.
@@ III.1 Cosmic string
Cosmic strings are topological defects similar to domain walls or monopoles, arising when the vacuum manifold $\mathcal{M}$ supports non-contractible loops. [^3] In more technical terms, the first homotopy group of the (true or false) vacuum manifold must be non-trivial, i.e. $\pi_{1}(\mathcal{M})\neq 0$. In the case of a complex scalar field with $U(1)$ symmetry, the field picks up a non-vanishing expectation value $|\phi|\simeq v_{f}/\sqrt{2}$ outside the string core, corresponding in Fig. 2 either to the global (dashed line) or the local (solid line) minimum. As mentioned before, in this work we are mostly interested in the latter, metastable case. As its defining property, the cosmic string has a non-vanishing winding number $n$ associated with its complex phase. Correspondingly, the (straight) cosmic string can be expressed as $\phi\simeq(v_{f}/\sqrt{2})e^{in\theta}$ sufficiently far away from the string core. Here, we adopted cylindrical coordinates $x^{\mu}=(t,r,\theta,z)$, where $\theta$ denotes the polar angle and $r=\sqrt{x^{2}+y^{2}}$ the radial distance from the string core at $r=0$. The $z$-axis is aligned with the string. Cosmic strings are topological solitons which cannot be continuously “un-winded”. In this work, we will assume that the string tension $\mu_{s}$ is small such that we can neglect its gravitational back-reaction. In particular, taking into account the string’s deficit angle of order $8\pi\,G\,\mu_{s}\ll 1$ would not alter the results of this work significantly.
@@
$A_{\theta}=\frac{n}{g}\,a(r)\quad\text{and}\quad A_{r}=A_{t}=A_{z}=0\,.$ (10)
@@
where $x=16\,n^{2}\epsilon$. The first inequality also guarantees $\Delta R/R\ll 1$, as assumed above. It is useful to distinguish two regimes. For $x\lesssim 1$, i.e. $\epsilon\lesssim 1/(16n^{2})$, the conditions reduce to $n^{2}\gg 0.21$. For $x\ll 1$, we instead require $\epsilon\ll 1/(16n^{2})$ and $n^{2}\gg 0.42$. Thus, our analysis is quantitatively reliable only for winding numbers $n\geq 3$, although we expect the qualitative results to extend to the unit winding case. We note that the above bounds on the winding $n$ can be rewritten as $\epsilon\ll 1$ (or equivalently as $\Delta V/V_{\mathrm{max}}\ll 1$), which is precisely the standard thin-wall requirement. Moreover, the precise bounds should be expected to change for potentials with different shapes.
@@ III.2 Bounce
where $\dot{R}=dR(\tau)/d\tau$. Geometrically, multiplying the radial distance by the $\gamma$-factor gives the (proper) distance from the wall surface in the four-dimensional Euclidean spacetime. It is nothing else than the Lorentz factor, which accounts for the fact that as the bubble wall moves with some velocity $\dot{R}=dR/d\tau$ its width gets contracted compared to a static wall with the same radius, which would be described as $f_{\mathrm{wall}}(|\mathbf{x}|-R)$. It is straightforward to show that the ansatz (43) satisfies (42) if we assume that $f_{\mathrm{wall}}^{\prime}\simeq f_{\mathrm{wall}}^{\prime\prime}\simeq 0$ outside the wall region and $\ddot{R}\,\Delta\mathcal{R}\ll 1$.
@@ V.1 Comparison of nucleation efficiency
$e^{2(b-1)\,B_{0}}\frac{1}{N_{s}^{2}b^{2}}\ll 1\,.$ (70)
Equivalently, for $b\lesssim 1/2$, string-induced vacuum decay proceeds more efficiently than the conventional bubble nucleation, while for $b\gtrsim 1/2$ the standard Coleman mechanism dominates. Using our numerical results from Eqs. (58) and (63), we find that this translates into the approximate parameter conditions
@@ V.3.3 String network
where $\Gamma^{(k)}=\Gamma k^{-n}/\sum_{p=1}^{\infty}p^{-n}$ is the fractional decay rate into the $k$-th harmonic with $n=4/3$ assuming cusps dominating the small-scale structure of the string [35], and $l_{k}=2k/\nu$ is the comoving length scale of the $k$-th harmonic of the string, corresponding to radiation observed at the frequency $\nu$ today. Moreover, $n_{\rm loop}(l,t)$ is the number density of strings of physical length $l$ at time $t$. It takes the form
@@
for $t_{\rm SSB}<t_{\rm ini}(l,t)<t<t_{\star}$ and is zero otherwise. Here $t_{\rm ini}(l,t)$ is the time of formation of the strings that, at time $t$, have physical size $l$. It is given by
@@
where $\alpha_{1}$, $\mathcal{F}$ and $C_{\rm eff}$ are all order unity constants. The later formula follows from
@@
In order to implement the vacuum decay of the string network at time $t_{\star}$, we introduce a hard cutoff in the time integration in (84). The standard result is then recovered for the choice $t_{\star}=t_{0}$. Assuming that the GW emission takes place deep inside the radiation era, $a(t)\propto\sqrt{t}$, and focusing on the first harmonic, the integral in (84) can be performed analytically,
@@
$\begin{split}\Omega_{\rm GW,0}^{\mathrm{(stable)}}&=\frac{\Gamma^{(1)}G\mu_{s}^{2}}{\rho_{0}}\frac{4\mathcal{F}C_{\rm eff}c^{4}}{3\alpha(\alpha+\Gamma G\mu_{s})}\frac{1}{{(\Gamma G\mu_{s})}^{3/2}}\\ &\sim 10^{-8}\sqrt{\frac{G\mu_{s}}{10^{-7}}}\,.\end{split}$ (89)
matches the usual expression for the flat GW spectrum emitted from a stable local string network. To arrive at the second line, we set $\Gamma=50$. The suppression of the spectrum caused by the vacuum decay is then captured by
@@
and assumed $\nu^{(2)}\gg\nu^{(1)}$, which holds for sufficiently sub-Planckian string tensions. There are three important frequency ranges: for $\nu\gg\nu^{(2)}$ the spectrum is approximately flat since $\mathcal{K}(\nu)\approx 1$, and one recovers the result for a stable network. In the intermediate range of frequencies $\nu^{(1)}<\nu<\nu^{(2)}$ one observes a tilted spectrum $\propto\mathcal{K}(\nu)\approx(\nu/\nu^{(2)})^{3/2}$. Around $\nu\approx\nu^{(1)}$ there is a sharper drop in the spectrum as $\mathcal{K}(\nu)\rightarrow 0$. Here one expects a causal $\propto\nu^{3}$ tail, as for any GW source that is localized in time. Such a spectrum is shown in Fig. 6 as the green line.
@@
Overall, the effect of the phase transition is to suppress the GW spectrum for frequencies below $\nu^{(2)}$, which is a consequence of limiting the time integration in (84) to times before $t_{*}$. Such a scenario might have interesting implications for bounds on local cosmic strings from PTAs, as well as the CMB. Stable local strings with $G\mu_{s}\gtrsim 10^{-10}$ are currently excluded by NANOGrav 15-year data [43]. However, as we demonstrate for the fiducial choice $G\mu_{s}=10^{-7}$ in Fig. 7, if the transition occurs at temperatures above $T_{*}\sim 50\,\rm eV$, the spectrum is suppressed around the frequencies probed by PTAs according to (92) and the bounds are evaded.
This mechanism of alleviating the bounds on local cosmic strings is somewhat similar to other metastable cosmic string scenarios. For example, in the scenario proposed in [6], strings can attach to magnetic monopoles and decay via monopole–antimonopole annihilation, suppressing the GW background [49, 50, 51, 60].
@@
Figure 7: GW spectra from a metastable local string network for different values of the temperature $T_{\mathrm{*}}$ of a symmetry restoring phase transition. The violin shapes indicate the uncertainties of the recent NANOGrav 15-year observations [44]. The chosen string tension, $G\,\mu_{s}=10^{-7}$ is typical of a GUT-scale transition and would be excluded for a stable network (red). If the phase transition occurs sufficiently early (yellow), the low-frequency part of the spectrum is suppressed, allowing the bound to be evaded. For an intermediate value, $T_{\mathrm{*}}\sim 50\,\mathrm{eV}$ (orange), corresponding to a post-BBN phase transition, the predicted and observed signals are of comparable magnitude.
@@
A more comprehensive analysis of the GW signatures from both local and global string networks in the context of string-induced vacuum decay is left for future work. For completeness, we briefly comment on networks of global strings. In contrast to local strings, global strings radiate predominantly into scalar waves corresponding to the massless Goldstone bosons of the spontaneously broken $U(1)$ symmetry. The resulting GW background is suppressed in amplitude relative to the local case and exhibits a tilted spectrum that decreases with frequency during the radiation era, rather than remaining flat [61].
@@ VI Conclusion
Beyond the possibility of evading GW constraints, it is intriguing to consider whether the observed PTA signal could in fact originate from a metastable cosmic-string network. As a first step in this direction, we have shown that a post-BBN vacuum phase transition can generate a spectrum with an amplitude consistent with current observations. We have also introduced a minimal two-field model that realizes the spontaneous symmetry breaking responsible for string formation and subsequently induces a delayed symmetry restoration through a string-induced vacuum decay. A more detailed investigation of this scenario is left for future work.
Finally, within the new early dark energy framework [66] it has recently been shown that a supercooled (hot new early dark energy[67, 68, 69, 70] or vacuum phase transition (cold new early dark energy [71, 72, 73]) occurring after BBN and injecting energy into the dark sector fluid can resolve the Hubble tension, a discrepancy in the observed value of the Hubble constant [74] (for a review see [75]). It is therefore natural to ask whether the phase transition discussed here is compatible with that framework and could provide a strong enough energy injection into the dark sector to address the tension while also explaining the NANOGrav signal (for an alternative idea see [76]). In particular, it might be possible to embed the the trigger mechanism discussed in [68] into the two-field model discussed in Sec. V.4.

### v2

@@ I Introduction
As mentioned above, one well-known decay mechanism of local cosmic strings involves the nucleation of monopole–antimonopole pairs along the string core [6], a process whose GW signatures were discussed for example in [49, 50, 51, 52, 53, 54] (see [43, 55, 56] for an analysis in the context of the latest NANOGrav data). In contrast, our scenario introduces a qualitatively new decay channel: instead of fragmenting through monopole pair production, the string network decays via $O(2)\times O(2)$ bubble nucleation. This process leads to unique features in the GW spectrum, making it potentially distinguishable from the monopole-induced scenario. While a complete comparative study of the GW signatures in different models is left for future work, here, we explore the possibility that the GW background observed by PTAs could arise from a string network that decays in a first-order phase transition. To that end, we fit the GW spectrum predicted by our model to the NANOGrav 15-year data.
Finally, we discuss a more comprehensive cosmological framework that explicitly includes the formation dynamics of the string network. The model features two interacting scalar fields, where a real scalar triggers the false vacuum decay of a complex scalar some time after the strings have formed.
@@ III.1 Cosmic string
Cosmic strings are topological defects, similar to domain walls or monopoles, that form when the vacuum manifold $\mathcal{M}$ admits non-contractible loops. [^3] More formally, this requires a non-trivial first homotopy group of the (true or false) vacuum manifold, $\pi_{1}(\mathcal{M})\neq 0$. In field theory, a canonical realization is provided by a complex scalar field with a spontaneously broken $U(1)$ symmetry [57]. Away from the string core, the field acquires a nonzero expectation value, $|\phi|\simeq v_{f}/\sqrt{2}$. In terms of the scalar field potential introduced above, this corresponds to either the global (dashed) or local (solid) minimum in Fig. 2. As mentioned before, in this work we are mostly interested in the latter, metastable case. As its defining property, the cosmic string has a non-vanishing winding number $n$ associated with its complex phase. Correspondingly, the (straight) cosmic string can be expressed as $\phi\simeq(v_{f}/\sqrt{2})e^{in\theta}$ sufficiently far away from the string core. Here, we adopted cylindrical coordinates $x^{\mu}=(t,r,\theta,z)$, where $\theta$ denotes the polar angle and $r=\sqrt{x^{2}+y^{2}}$ the radial distance from the string core at $r=0$. The $z$-axis is aligned with the string. Cosmic strings are topological solitons which cannot be continuously “un-winded”. In this work, we will assume that the string tension $\mu_{s}$ is small such that we can neglect its gravitational back-reaction. In particular, taking into account the string’s deficit angle of order $8\pi\,G\,\mu_{s}\ll 1$ would not alter the results of this work significantly.
@@
$A_{\theta}(x)=\frac{n}{g}\,a(r)\quad\text{and}\quad A_{r}=A_{t}=A_{z}=0\,.$ (10)
@@
where $x=16\,n^{2}\epsilon$. The first inequality also guarantees $\Delta R/R\ll 1$, as assumed above. It is useful to distinguish two regimes. For $x\lesssim 1$, i.e. $\epsilon\lesssim 1/(16n^{2})$, the conditions reduce to $n^{2}\gg 0.21$. For $x\ll 1$, we instead require $\epsilon\ll 1/(16n^{2})$ and $n^{2}\gg 0.42$. Thus, our analysis is quantitatively reliable only for winding numbers $n\geq 3$, although we expect the qualitative results to extend to the unit winding case. This would suggest our analysis is only valid for large enough winding number; however, we show in Fig. 3 that even for unit winding the thin-wall profile (18) (dashed) provides a good description of the full numerical solution (solid). Here, the numerical solutions are obtained via the gradient flow method, where we promote $f(r)\rightarrow f(r,s)$ and ‘evolve’ the profile according to
$\frac{\partial f}{\partial s}=f^{\prime\prime}+\frac{1}{r}f^{\prime}-\frac{n^{2}}{r^{2}}f-\frac{dV}{df},$ (27)
subject to the appropriate boundary conditions and starting from some guessed initial configuration. As the evolution parameter $s$ increases, the configuration relaxes toward a ‘static’ solution, which by construction corresponds to the correct string profile.
We note that the above bounds on the winding $n$ can be rewritten as $\epsilon\ll 1$ (or equivalently as $\Delta V/V_{\mathrm{max}}\ll 1$), which is precisely the standard thin-wall requirement. Moreover, the precise bounds should be expected to change for potentials with different shapes.
Figure 3: Field profiles of global strings for the potential in (6) with $\epsilon=0.005$. We compare the full numerical solution, obtained using the gradient flow method (solid lines), with the thin-wall approximation (dashed lines) and find good agreement.
@@ III.2 Bounce
where $\dot{R}=dR(\tau)/d\tau$. Geometrically, multiplying the radial distance by the $\gamma$-factor gives the (proper) distance from the wall surface in the four-dimensional Euclidean spacetime. It is nothing else than the Lorentz factor, which accounts for the fact that as the bubble wall moves with some velocity $\dot{R}=dR/d\tau$ its width gets contracted compared to a static wall with the same radius, which would be described as $f_{\mathrm{wall}}(|\mathbf{x}|-R)$. It is straightforward to show that the ansatz (44) satisfies (43) if we assume that $f_{\mathrm{wall}}^{\prime}\simeq f_{\mathrm{wall}}^{\prime\prime}\simeq 0$ outside the wall region and $\ddot{R}\,\Delta\mathcal{R}\ll 1$. As a consistency check, we will show a posteriori that the last relation is indeed satisfied.
@@ V.1 Comparison of nucleation efficiency
$e^{(2b-1)\,B_{0}}\frac{1}{N_{s}^{2}b^{2}}\ll 1\,.$ (71)
Equivalently, assuming $N_{s}$ of order unity, for $b\lesssim 1/2$, string-induced vacuum decay proceeds more efficiently than the conventional bubble nucleation, while for $b\gtrsim 1/2$ the standard Coleman mechanism dominates. Using our numerical results from Eqs. (59) and (64), we find that this translates into the approximate parameter conditions
@@ V.3.3 String network
where $\rho_{0}$ is the total energy density today, i.e. $\rho_{0}=\rho_{\mathrm{tot}}(t_{0})$, and $\Gamma^{(k)}=\Gamma k^{-n}/\sum_{p=1}^{\infty}p^{-n}$ is the fractional decay rate into the $k$-th harmonic with $n=4/3$ assuming cusps dominating the small-scale structure of the string [35], and $l_{k}=2k/\nu$ is the comoving length scale of the $k$-th harmonic of the string, corresponding to radiation observed at the frequency $\nu$ today. Moreover, $n_{\rm loop}(l,t)$ is the number density of strings of physical length $l$ at time $t$. In the scaling regime it takes the form
@@
for $t_{\rm SSB}<t_{\rm ini}(l,t)<t<t_{0}$ and is zero otherwise. Here $t_{\rm ini}(l,t)$ is the time of formation of the strings that, at time $t$, have physical size $l$. It is given by
@@
where $\alpha_{1}\approx 0.1$, $\mathcal{F}\approx 0.1$ and $C_{\rm eff}\approx 5.3(0.4)$ during the radiation (matter)-dominated era. The later formula follows from
@@
In order to implement the vacuum decay of the string network, we introduce a cutoff at $t=t_{\star}$ in the time integration in (85). This is equivalent to setting $n_{\rm loop}(l,t)=0$ for $t>t_{\star}$. The standard result is then recovered for the choice $t_{\star}=t_{0}$. Assuming that the GW emission takes place deep inside the radiation era, $a(t)=c\sqrt{t}$, and focusing on the first harmonic with $k=1$, the integral in (85) can be performed analytically,
@@
$\begin{split}\Omega_{\rm GW,0}^{\mathrm{(stable)}}&=\frac{\Gamma^{(1)}G\mu_{s}^{2}}{\rho_{0}}\frac{4\mathcal{F}C_{\rm eff}c^{4}}{3\alpha}\frac{(\alpha+\Gamma G\mu_{s})^{3/2}}{{(\Gamma G\mu_{s})}^{3/2}}\\ &\sim 4\times 10^{-9}\sqrt{\frac{G\mu_{s}}{10^{-7}}}\,.\end{split}$ (90)
matches the usual expression for the flat GW spectrum emitted from a stable local string network. To arrive at the second line, we set $\Gamma=50$ and $c=1.1(H_{0}^{2}\Omega_{\mathrm{rad},0})^{1/4}$. The suppression of the spectrum caused by the vacuum decay is then captured by
@@
where $\nu^{(2)}\gg\nu^{(1)}$ for sufficiently sub-Planckian string tensions. There are three important frequency ranges: for $\nu\gg\nu^{(2)}$ the spectrum is approximately flat since $\mathcal{K}(\nu)\approx 1$, and one recovers the result for a stable network. In the intermediate range of frequencies $\nu^{(1)}<\nu<\nu^{(2)}$ one observes a tilted spectrum $\propto\mathcal{K}(\nu)\approx(\nu/\nu^{(2)})^{3/2}$. Around $\nu\approx\nu^{(1)}$ there is a sharper drop in the spectrum as $\mathcal{K}(\nu)\rightarrow 0$. Here one expects a causal $\propto\nu^{3}$ tail, as for any GW source that is localized in time. Such a spectrum is shown in Fig. 7 as the green line.
@@
Overall, the effect of the phase transition is to suppress the GW spectrum for frequencies below $\nu^{(2)}$ according to (93), which is a consequence of limiting the time integration in (85) to times before $t_{*}$. Such a scenario can have interesting implications for bounds on local cosmic strings from PTAs, as well as the CMB. While stable local strings with $G\mu_{s}\gtrsim 10^{-10}$ are currently excluded by NANOGrav 15-year data [43], below we demonstrate how these bounds can be evaded in the presence of a symmetry-restoring phase transition.
[…]
@@
Figure 8: Left: GW spectra from a metastable local string network for different values of the temperature $T_{\mathrm{*}}$ of a symmetry restoring phase transition. The violin shapes indicate the uncertainties of the recent NANOGrav 15-year observations [44]. The chosen string tension, $G\,\mu_{s}=10^{-7}$ is typical of a GUT-scale transition and would be excluded for a stable network (red). If the phase transition occurs sufficiently early (yellow), the low-frequency part of the spectrum is suppressed, allowing the bound to be evaded. For an intermediate value, $T_{\mathrm{*}}=170\,\mathrm{eV}$ (orange), corresponding to a post-BBN phase transition, the predicted and observed signals are of comparable magnitude. Right: Reconstructed $1\sigma$ and $2\sigma$ confidence intervals from a frequentist analysis of the NANOGrav 15-year data. The values of $G\,\mu_{s}$ and $T_{\mathrm{*}}$ corresponding to the best fit and to the orange and yellow lines in the left panel are shown as black, orange and yellow dots, respectively. Overall, our scenario provides an excellent fit to NANOGrav data with a best fit corresponding to $\chi^{2}/\text{dof}\approx 0.5$.
In the right panel of Fig. 8 we show the results of a frequentist analysis of the NANOGrav 15-year data, following the method in [63]. We approximate the posterior distributions in each frequency bin, i.e. the gray ‘violins’, as piecewise Gaussian in $\log_{10}\Omega_{\mathrm{GW}}$, using the median $\log_{10}\bar{\Omega}_{\mathrm{GW},i}$ as the central value and separate variances $\Delta\log_{10}{\Omega}_{\mathrm{GW},i}$ for upward and downward fluctuations. For given values of $\mu_{s}$ and $T_{\star}$, we compute $\chi^{2}$ as
$\chi^{2}=\sum_{i\in\rm bins}\frac{(\log_{10}\Omega_{\rm GW}(\nu_{i})-\log_{10}\bar{\Omega}_{\mathrm{GW},i})^{2}}{(\Delta\log_{10}{\Omega}_{\mathrm{GW},i})^{2}},$ (94)
summing only over the first $14$ bins to eliminate the high-frequency noise. The best-fit parameters (shown as the black dot) correspond to $\chi^{2}_{\rm min}$, and the $1\sigma$ ($2\sigma$) contours are identified with $\chi^{2}-\chi^{2}_{\rm min}=2.3\,(6.2)$. The region above the contours evades the PTA bounds. Inside the contours the predicted GW signal itself provides a good fit to data. In particular, we find $\chi^{2}/\text{dof}\approx 0.5$ for the best fit (shown as black dot) and $\chi^{2}/\text{dof}\approx 0.8$ for parameters corresponding to the orange line in the left panel (shown as orange dot in the right panel). The number of degrees of freedom (dof) is 12=14212=14-2, where we subtract the number of model parameters. In contrast, the best fit value for the case of stable strings is $\log_{10}(G\mu_{s})\approx-10.2$ and has $\chi^{2}/\text{dof}\approx 1.7$. An abrupt increase of $\chi^{2}$ for $G\mu_{s}>10^{-4}$ is due to the pulsars becoming sensitive to frequencies around $\nu^{(1)}$, where the spectrum has a sharp drop. This abrupt behavior is an artifact of describing the symmetry-restoring phase transition as an instantaneous process.
In summary, PTA data prefers a symmetry-restoring phase transition occurring between BBN and recombination. Such a transition is not constrained by the CMB provided the latent heat is small. This is achieved when the two minima of the potential are nearly degenerate, so that the vacuum energy difference is small compared to the radiation energy density. Moreover, the string network carries only a negligible fraction $\sim G\mu_{s}$ of the total energy density, so its decay does not lead to observable entropy production or distort the CMB. Consequently, such a late transition can occur without conflicting with any cosmological bounds.
This mechanism of alleviating the bounds on local cosmic strings is similar to other metastable cosmic string scenarios. For example, in the scenario proposed in [6], strings can attach to magnetic monopoles and decay via monopole–antimonopole annihilation, suppressing the GW background [49, 50, 51, 64].
@@
For completeness, we briefly comment on networks of global strings. In contrast to local strings, global strings radiate predominantly into scalar waves corresponding to the massless Goldstone bosons of the spontaneously broken $U(1)$ symmetry. The resulting GW background is suppressed in amplitude relative to the local case and exhibits a tilted spectrum that decreases with frequency during the radiation era, rather than remaining flat [65]. A more comprehensive analysis of the GW signatures from both local and global string networks in the context of string-induced vacuum decay is left for future work.
@@ VI Conclusion
Beyond the possibility of evading GW constraints, it is intriguing to consider whether the observed PTA signal could in fact originate from a metastable cosmic-string network. As a first step in this direction, we have shown that a post-BBN vacuum phase transition can generate a spectrum with an amplitude that provides a good fit to NANOGrav data. We have also introduced a minimal two-field model that realizes the spontaneous symmetry breaking responsible for string formation and subsequently induces a delayed symmetry restoration through a string-induced vacuum decay. A more detailed investigation, which compares different metastable string scenarios and uses the data fit to constrain the details of the two-field model, is left for future work.
Finally, within the new early dark energy framework [70] it has recently been shown that a supercooled (hot new early dark energy[71, 72, 73, 74] or vacuum phase transition (cold new early dark energy [75, 76, 77]) occurring after BBN and injecting energy into the dark sector fluid can resolve the Hubble tension, a discrepancy in the observed value of the Hubble constant [78] (for a review see [79]). It is therefore natural to ask whether the phase transition discussed here is compatible with that framework and could provide a strong enough energy injection into the dark sector to address the tension while also explaining the NANOGrav signal (for an alternative idea see [80]). In particular, it might be possible to embed the trigger mechanism discussed in [72] into the two-field model discussed in Sec. V.4.